The minute hand on a watch is 8 mm long and the hour hand is 6 mm long. How fast is the distance between the tips of the hands changing at one o'clock? (Round your answer to one decimal place.)
Answer #1

The rough sketch for hour hand and minute hand is shown as below:

6 8

The hour hand trawels around the clock at |revoluti on 2 radian π 12h 12h The minute hand travels around the clock at |revolu

Put θ--in equation (3)s 兀 D-100-96 cos- → D = 100-96 Now, substitute D-100 dt= 6_ , g=-,In equation (4 ) dD 48sin 6 dte dt D